The horse racing record for a 1.50 mi. track is shared by two horses: Fiddle Isle, who ran the race in 143 s on March 21, 1970, and John Henry, who ran the same distance in an equal time on March 16, 1980. What were the horses' average speeds in mi/s
1
Expert's answer
2020-10-19T13:20:05-0400
"v=\frac{s}{t}=\frac{1.5}{143}=0.0105\frac{mi}{s}"